For the given null ( ?0 ) and alternative ( ?1 ) hypotheses, find the positive value, ? , and the negative value, −? , that define the critical region (rejection region) for a one-sample two-tailed ? ‑test for a mean with a significance level of ?=0.040 . ?0:?=5,?1:?≠5. Use software to find the values of ? to at least three decimal places. You may find one of these software manuals useful. −?= ? ?=?
Answer #1

Solution :

This is the two tailed test,  

The null and alternative hypothesis is ,

H0 :  \mu = 5

Ha : \mu   \neq 5

\alpha = 0.040

\alpha/2 = 0.020

Z\alpha/2 = Z0.020 = 2.054

-z < -2.054 and z > 2.054